OpenStudy (anonymous):

whats the square root of 126

OpenStudy (anonymous):

11.225( correct to 3 dp)

OpenStudy (anonymous):

the calculator will tell u

OpenStudy (anonymous):

i need it no in a decimal

OpenStudy (anonymous):

then its approx 11

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{126}\] \[\large \sqrt{9*14}\] \[\large \sqrt{9}\sqrt{14}\] \[\large 3\sqrt{14}\] So \[\large \sqrt{126}=3\sqrt{14}\]

OpenStudy (anonymous):

126 = 9 * 14 sqrt 126 = 3sqrt 14

OpenStudy (anonymous):

then it is not a complete solution u have jst simplified it

OpenStudy (anonymous):

they are asking solution !!!!